# Beyond Exact Match: Semantically Reassessing Event Extraction by Large Language Models

### v1

@@ 3.3.3 Ablation of Adaptive Mechanism
Moreover, it can also be observed that the adaptive mechanism significantly improves the percent agreement and Spearman correlation between RAEE and human average in most cases, i.e., over 2.85% percent agreement and 26.34 Spearman correlation absolute improvements. This is because pre-defined criteria help reduce the randomness in the LLM evaluation.
Summary
While LLM slightly underperforms human average, it significantly outperforms EM and achieves a strong correlation with human judgments. It is important to note that RAEE is a trade-off solution by considering the evaluation cost and reliability, which could reliably assess the performance of event extraction results with affordable cost. For example, as shown in Table 2, the average time and financial cost of GPT-4o on each evaluated model and each dataset are much smaller than those of pure human annotation, proving the scalability and reliability of LLM judge models in our proposed RAEE framework.
Evaluation Cost Time (hours) Financial (\$)
RAEE + GPT-4o $\approx 0.23$ $\leq 5$
RAEE + Human $\approx 19.22$ $\approx 182.32$
Table 2: The average evaluation cost of human annotation and GPT-4o for each evaluated model on each dataset.
@@
4 Re-Assessment Experiments
The reliability of RAEE with GPT-4o as the judging agent has been proven. In this section, we reassess the performance of existing event extraction models in numerous datasets across various domains by using our proposed RAEE evaluation framework.
@@ 4.1 Experimental Settings
ED and EAE event extraction tasks are evaluated respectively. Following Lin et al. (2020), we focus on trigger classification (TC) and argument classification (AC). The semantic-level f1-score for TC and AC is described in Section 3.1, using specialized prompts to assess both precision and recall.
@@
We reassess the overall 8 state-of-the-art event extraction models with different frameworks (5 models for ED and EAE tasks). These models can be broadly classified into two categories: (1) Discriminative models that classify each token of text into predefined categories, including Tagprime-C Hsu et al. (2023a), OneIE Lin et al. (2020), Query&Extract Wang et al. (2022), CEDAR Li et al. (2023a); (2) Generative models that produce event-related outputs directly, including BartGen Li et al. (2021), PAIE Ma et al. (2022), DEGREE Hsu et al. (2022), AMPERE Hsu et al. (2023b). All models train and inference under the unified TextEE framework Huang et al. (2024), with implementation details shown in Appendix E.
@@
We use an overall 10 datasets to test the performance of models, covering diverse domains, including news of ACE05-en Doddington et al. (2004) and RAMS Ebner et al. (2020), general domain of FewEvent Deng et al. (2020) and GENEVA Parekh et al. (2023), Wikipedia of MEE-en Pouran Ben Veyseh et al. (2022) and WikiEvents Li et al. (2021), biomedical domain of Genia2011 Kim et al. (2011), multimedia of M2E2 Li et al. (2020), cybersecurity of CASIE Satyapanich et al. (2020), and Pharmacovigilance of PHEE Sun et al. (2022). The statistics of these datasets are provided in Appendix B.
@@ 4.2 Main Results
Table 3 and 4 show EM evaluations and RAEE reassessment results for ED and EAE tasks, respectively. Notably, across all models, RAEE reassessment shows a significant improvement in metrics compared to EM evaluations. On average, the precision of all models increased by more than 20%, with recall improvement exceeding 10% for ED and 15% for EAE. This indicates that EM evaluation severely underestimates model performance by overlooking a substantial number of semantically correct predictions, while RAEE provides a more accurate evaluation of model performance across diverse datasets.
@@
Table 5: Average f1-score difference between EM and RAEE evaluation. It can be found that generative models are significantly more underestimated compared to discriminative models.
Moreover, as shown in Table 5, compared to discriminative models, the performance gap between RAEE and EM is significantly larger for generative models, indicating that EM introduces a greater bias for generative models.
Furthermore, the rank of models across different datasets changes considerably under RAEE. For instance, in the ED task, DEGREE ranks lower than OneIE and CEDAR on average under EM evaluation, while DEGREE outperforms these two models under RAEE. This suggests that EM evaluation distorts the true performance rank of models, leading to misconceptions about the model performance.
@@ 4.3 Fine-grained Reason Analysis
We further conduct a fine-grained analysis to uncover the main reasons behind the misjudgments by EM, as well as the primary failure modes of existing event extraction models under our proposed RAEE evaluation framework. Specifically, 200 sampled data items for both ED and EAE tasks are annotated through multiple human experts.
Reasons of EM Misjudgements
As for instances evaluated as incorrect under EM but reassessed as correct under RAEE, we summarize 4 main categories of reasons for ED and EAE tasks, through human annotations: (1) With CoreWord: Compared to the ground-truth, the predicted results contain the core words and only have missing or redundant words that do not affect the overall semantics; (2) Co-reference: There is a relationship between the predicted result and a certain ground-truth refer to the same entity or concept; (3) Correct: While no match is found in ground-truth, these instances are still reassessed as correct; (4) Other non-representative reasons.
@@
As shown in Figure 3, it can be observed that the most misjudgments for EM evaluation in ED and EAE tasks are CoreWord and Correct. This indicates that current event extraction models tend to retain core words while expanding or omitting some less relevant words without changing the semantic meaning of the extracted results. Additionally, this also suggests that the annotations of existing datasets may not be comprehensive, as there are still a significant number of extracted results that are reasonable and correct but are not included in the ground-truth.
Failure Modes of Existing Models Under RAEE
As for instances evaluated as incorrect under both EM and RAEE, we also summarize 4 kinds of reasons for ED and EAE tasks, through human annotations: (1) Without CoreWord: Compared to the ground-truth, the predicted results have missing core words that convey the primary semantics; (2) WrongType: Even though there is a semantic-level match of span in the ground-truth, the predicted results have an incorrect type of event or role; (3) Incorrect: These instances are reassessed as incorrect and they are not included in the ground-truth; (4) Other non-representative reasons.
Figure 4: Distribution of failure modes under our proposed RAEE evaluation framework on ED and EAE tasks.
As shown in Figure 4, the most frequent failure modes of existing models on ED and EAE tasks are WrongType. This phenomenon indicates that although existing models could locate the spans correctly, their types are usually misclassified.This suggests that future works should focus on improving the capabilities of models in understanding and classifying the types of triggers and arguments.
@@
5 Do LLMs Solve EE Task?
Under EM evaluation system, LLMs have a significant performance gap with state-of-the-art small fine-tuned models on EE task Huang et al. (2024). In this section, we also reassess the capabilities of LLMs. Specifically, 6 advanced open-source and closed-source LLMs are included: GPT-4o, GPT-3.5-turbo, Claude-3-Opus, Llama-3.1-70B-Chat, Qwen2-70B-Chat Yang et al. (2024) and Mixtral-8X7B-instruct Jiang et al. (2024). The well-selected prompts for inference on ED and EAE tasks are shown in Appendix D.2.
Figure 6 and 7 demonstrate results of LLMs and Top-2 fine-tuned small models under EM and RAEE evaluation on ED and EAE tasks, respectively. It can be found that LLMs generally underperform compared to fine-tuned models under EM evaluation, with significantly lower f1-scores on ED and EAE tasks. However, LLMs show remarkable improvements under RAEE evaluation, with f1-scores much closer to those of fine-tuned models, indicating that the event extraction capabilities of LLMs have been greatly underestimated. Notably, f1-score improvement of LLMs on EAE task, over 50%, is higher than that on ED task, over 30%. Our human annotations over hundreds of cases reveal that LLMs are more likely to generate out of pre-defined event type labels in ED task than role type labels in EAE task, as reflected in their relatively lower recall scores of ED.
Additionally, we also conduct few-shot experiments with the best-performing LLM, GPT-4o [^2] https://openai.com/index/hello-gpt-4o/. It is found that as the number of few-shot examples increases, the performance of the model fluctuates on ED task and declines on EAE task. Our human annotations reveal that these results are heavily influenced by the quality of selected shots. With data to be extracted and sampled from datasets in a diverse range of domains, it is more likely to mislead the model by providing a few shots from only a few specific domains. We will leave a more detailed analysis of few-shot prompting strategies in LLMs for our future work.
@@ 6 Conclusion and Future Works
In this paper, we propose a reliable automatic evaluation framework, RAEE, for assessing the event extraction task at the semantic level rather than the token level of exact match evaluation. Extensive experimental results demonstrate that GPT-4o achieves a strong correlation with human judgments within our RAEE framework, establishing RAEE as a reliable automatic evaluation framework for assessing the quality of event extraction results. Additionally, we reassess the performance of numerous prior works and advanced LLMs in the event extraction task using our proposed RAEE framework. Extensive experimental results indicate that the capabilities of prior works and LLMs have been significantly underestimated. Finally, the fine-grained reason analysis reveals several intriguing phenomena worth further exploration.
In the future, we will conduct a more comprehensive evaluation for event extraction tasks to reassess the performance of existing models. Besides, we will extend the evaluation of model performance with RAEE from closed domains to the open domain.
Appendix A Limitations
A.1 LLM Judgments v.s. Human Judgments
In this paper, advanced LLM GPT-4 is used to automatically judge the correctness of the predicted trigger words and arguments for the event extraction task. However, it should be noted that LLMs cannot always perfectly align with human judgments, which is the most reliable evaluation. For example, GPT-4o achieves comparable agreement and correlations with human average performance in ED and EAE annotations. Since the ED precision task is challenging, the Spearman correlation score on it is significantly lower than human performance. While we have made efforts to improve the relevance between LLM and human assessment through prompt design, this gap still exists. In the future, we would like to discover other strategies to further improve the reliability of LLM-as-a-Judge, such as multi-agent debate Chan et al. (2023).
A.2 Computational Cost of Evaluating One Model
Unlike existing EM evaluation, our proposed metrics rely on advanced LLMs to annotate the quality of predicted events, leading to much higher costs. The average cost of evaluating one model on each benchmark is between 2-5\$, which is smaller than established benchmarks like AlpacaEval Li et al. (2023b).
A.3 Insufficient Evaluation for LLMs
@@
A.4 Other Advanced LLMs as Evaluator
In this paper, the GPT-4 model is used to conduct the automatic evaluation in RAEE. Due to the budget limitation, we did not employ a diverse set of advanced LLMs for this task, such as Claude-3 and PaLM. Since our preliminary study reveals that GPT-4o is slightly better than other advanced LLMs, we mainly use it as the automatic judge model in our proposed RAEE framework.
Appendix B Datasets Statistics
We utilize 10 widely-used datasets as evaluation benchmarks in both EM and RAEE evaluation. Since these datasets come from different domains and vary in pre-defined ontologies and size, it is necessary to introduce basic statistics of these datasets, as shown in Table 8.
@@ Appendix B Datasets Statistics
Appendix C Key Criteria in RAEE
Considering that the requirements for EE vary across different scenarios and applications, we introduce key criteria in RAEE to enable the adaptive mechanism. As shown in Tabel 9, these key criteria are based on factors that significantly influence the evaluation results in our human annotation meta-evaluation experiment. Furthermore, this adaptive mechanism is easily extendable, allowing for the addition of personalized judgment criteria to account for more factors in extraction.
@@ Appendix C Key Criteria in RAEE
Appendix D Prompts
D.1 Reassessment Prompts in RAEE
@@
D.2 Inference Prompts for LLMs Extracting Events
Table 12 and 14 illustrate the prompts we use for testing the event extraction ability of LLMs in ED and EAE tasks, respectively.
@@ D.2 Inference Prompts for LLMs Extracting Events
Appendix E Implementation Details

### v2

@@ 3.3.2 Ablation of Adaptive Mechanism
The proposed adaptive mechanism improves correlation with human in most cases, regardless of the evaluation agent. Besides, except for the deepseek series LLMs [^4] The technical report of deepseek mentions that its LLM is sensitive to prompts DeepSeek-AI et al. (2025). , the adaptive mechanism significantly enhances the reliability of RAEE, as evidenced by a notable reduction in standard deviations across most evaluation results. For example, the standard deviation for GPT-4o decreased by at least 0.36 (EAE precision) and by up to 4.46 (EAE recall) across all tasks. This validates the effectiveness of the adaptive mechanism.
3.3.3 Impart of Evaluation Agents Selection
It is observed that the evaluation quality of RAEE depends on the capability of the evaluation agent. Recently, deepseek-r1 has been validated as the most comprehensive model across various tasks DeepSeek-AI et al. (2025). In RAEE, deepseek-r1 delivers the overall strongest evaluation performance. Notably, it achieves a very strong correlation with human judgments in the EAE recall task, with a Percent agreement of 93.49% and a Spearman correlation of 0.87, significantly outperforming other LLMs.
It is important to note that RAEE is a trade-off solution by balancing evaluation cost and reliability. When traditional methods fail to provide reliable semantic-level evaluation and pure human evaluation is costly, RAEE offers a relatively reliable and affordable alternative. As shown in Table 2, the average time and financial cost of RAEE are significantly lower than those of pure human annotation. Furthermore, RAEE is not restricted by the choice of specific evaluation agents. With the improvement in the text evaluation capabilities of LLMs, the performance of RAEE will also improve, further reinforcing its reliability.
Evaluation Methods Time (hours) Financial (\$)
RAEE + Deepseek-r1 $\approx 5.85$ $\approx 5.44$
RAEE + GPT-4o $\approx 0.20$ $\approx 17.79$
Pure Human $\approx 29.27$ $\approx 255.13$
Table 2: The average evaluation cost of human annotation and RAEE for an evaluated model on a dataset (approximate 1170 pieces of data on average). Assume RAEE evaluates with 10 parallel programs.
@@ 3.3.3 Impart of Evaluation Agents Selection
4 Reassessment Experiments
The reliability of RAEE has been validated. In this section, we will apply RAEE to reassess the performance of existing EE models in mainstream datasets across various domains.
@@ 4.1 Experimental Settings
ED and EAE tasks are evaluated separately. Following Lin et al. (2020), we focus on trigger classification (TC) and argument classification (AC) tasks. The semantic-level f1-score for TC and AC is described in Section 3.1. Due to the time constraints, we select GPT-4o as the evaluation agent for RAEE.
@@
We reassess 8 state-of-the-art event extraction models using different frameworks. These models can be broadly categorized into two types: (1) Discriminative models that classify tokens into predefined categories, including Tagprime-C Hsu et al. (2023a), OneIE Lin et al. (2020), Query&Extract Wang et al. (2022), and CEDAR Li et al. (2023a); (2) Generative models that directly produce event-related outputs, including BartGen Li et al. (2021), PAIE Ma et al. (2022), DEGREE Hsu et al. (2022), and AMPERE Hsu et al. (2023b). Implementation details are shown in Appendix D.
@@
We evaluate the performance of event models on 10 datasets covering diverse domains, including news ACE05-en Doddington et al. (2004) and RAMS Ebner et al. (2020), general FewEvent Deng et al. (2020) and GENEVA Parekh et al. (2023), Wikipedia MEE-en Pouran Ben Veyseh et al. (2022) and WikiEvents Li et al. (2021), biomedical Genia2011 Kim et al. (2011), multimedia M2E2 Li et al. (2020), cybersecurity CASIE Satyapanich et al. (2020), and pharmacovigilance PHEE Sun et al. (2022). Dataset statistics are provided in Appendix A.
@@ 4.2 Main Results
Table 3 and 4 show both EM evaluations and RAEE reassessment results for ED and EAE tasks, respectively. Notably, across all models, RAEE reassessment shows a significant improvement in metrics compared to EM. On average, the precision of all models increased by more than 20.52%, with recall improvement exceeding 9.85% for ED and 16.71% for EAE. It indicates that EM severely underestimates model performance by overlooking a substantial number of semantically correct predictions. Moreover, as shown in Table 5, the evaluated models’ performance gap between RAEE and EM is significantly larger for generative models compared to discriminative models, indicating that EM introduces a greater bias for generative models.
@@
Table 5: Average f1-score gap between EM and RAEE.
Furthermore, in many datasets, model rankings change considerably when shifting EM to RAEE. For instance, in the ED task, DEGREE ranks lower than OneIE and CEDAR across all datasets under EM, but outperforms both on average under RAEE. In the EAE task, rankings shift in ACE05, Genia2011 and CASIE datasets. It suggests that EM distorts the true performance ranks of models, leading to misconceptions about model performance.
@@ 4.3 Fine-grained Reason Analysis
To explore the main reasons behind the misjudgments evaluated by EM and the primary failure modes in RAEE, we conduct a fine-grained analysis experiment, extending the reasons reported in Han et al. (2024); Chen et al. (2024a). 200 sampled pieces of data for both ED and EAE tasks are evaluated through multiple human experts.
Misjudgments under EM
As for instances evaluated as incorrect under EM but reassessed as correct under RAEE, we summarize 4 main categories of reasons: (1) With CoreWord: Compared to the ground-truth, the predicted results contain the core words and only have missing or redundant words that do not affect the overall semantics; (2) Co-reference: There is a relationship between the predicted result and a certain ground-truth refer to the same entity or concept; (3) Unannotated Correct: While no match is found in ground-truth, the predicted results are still reassessed as correct; (4) Other non-representative reasons.
@@
As shown in Figure 3, the main misjudgments in EM are CoreWord and U-Correct. It indicates that current EE models tend to preserve core words while expanding or omitting some less relevant words, without changing the semantic meaning of extractions. The proportion of U-Correct is similar to the 38.70% reported in Han et al. (2024). Two real examples are provided in Figure 4. This suggests that it may be necessary to introduce new rules, such as supplementary annotations, to enhance the comprehensiveness of some datasets.
Figure 4: Two examples of Unannotated Correct cases.
Failure Modes under RAEE
As for instances evaluated as incorrect under both EM and RAEE, we also summarize 4 kinds of reasons: (1) Without CoreWord: Compared to the ground-truth, the predicted results have missing core words that convey the primary semantics; (2) WrongType: Even though there is a semantic-level match of span in the ground-truth, the predicted results have an incorrect type of event or role; (3) Unannotated Incorrect: These instances are reassessed as incorrect and they are not included in the ground-truth; (4) Other non-representative reasons.
Figure 5: Distribution of failure modes under our proposed RAEE evaluation framework on ED and EAE.
As shown in Figure 5, the most frequent failure mode is WrongType. It indicates that while existing EE models correctly locate the spans, they often misclassify the types. Future works should focus on enhancing models’ ability to accurately understand and classify the types of extractions.
5 LLMs Performance in EE Task
Previous work has validated a significant performance gap between LLMs and state-of-the-art smaller fine-tuned models in the EE task under EM Huang et al. (2024). In this section, we try to reassess LLMs performance in the EE task. Specifically, 6 advanced open-source and closed-source LLMs are reassessed: GPT-4o, GPT-3.5-turbo, Claude-3-Opus, Llama-3.1-70B, Qwen2-70B Yang et al. (2024) and Mixtral-8X7B Jiang et al. (2024). The well-selected prompts for inference are shown in Appendix C.2.
Experiment results are shown in Table  6 and 7. It is found that LLMs generally underperform smaller fine-tuned models in both ED and EAE task, regardless of evaluation. However, LLMs show much more improvements than smaller models when shifting EM to RAEE, indicating that LLMs’ event extraction performance is greatly underestimated. Notably, f1-score improvement of LLMs on EAE task, over 51.92%, is higher than that on ED task, over 30.74%. This is mainly due to the fact that extractions in EAE task consist of more words (2.52 on average) than those in ED task (1.21 on average), making them more susceptible to misjudgment by EM.
@@ 5 LLMs Performance in EE Task
We also conduct few-shot experiments with the best-performing LLM, GPT-4o. It is found that as the number of few-shot examples increases, the model’s performance fluctuates on ED task and declines on EAE task. Our human annotations reveal that these results are heavily influenced by the quality of selected shots. Since the data to be extracted is sampled from datasets across diverse domains, using few-shot examples from only a limited set of domains is more likely to mislead the model.
@@ 6 Conclusion and Future Works
In this paper, we introduce RAEE, a reliable semantic-level evaluation framework for event extraction, which moves beyond the token-level exact match evaluation. Extensive experiments validate RAEE’s strong correlation with human judgments, confirming its reliability. We apply RAEE to reassess the performance of numerous prior works and advanced LLMs in EE. Extensive experiments indicate that both prior works and LLMs are significantly underestimated. A further fine-grained reason analysis reveals several intriguing phenomena worth further exploration.
In the future, we will extend the evaluation of model performance with RAEE from closed domains to the open domain.
Limitations
LLM Judgments v.s. Human Judgments
In this paper, the advanced LLM GPT-4o is used to automatically judge predictions from event extraction models. However, it should be noted that LLMs cannot always perfectly align with human judgments, which is the most reliable evaluation. For example, GPT-4o achieves comparable agreement and correlations with human average performance in ED and EAE evaluations. While we have made efforts to improve the relevance between LLM and human assessment through prompt design, this gap still exists. In the future, we would like to discover other strategies to further improve the reliability of LLM-as-a-Judge, such as multi-agent debate Chan et al. (2023).
Sensitivity of LLMs to Prompts
In our preliminary study, it is found that the evaluation agent LLMs are sensitive to prompts. In the meta-evaluation experiments, using different prompts significantly affects LLMs’ correlation with human judgments. To mitigate this problem, we provide well-designed prompts that performs well across all LLMs in the meta-evaluation experiments, which can be found in Appendix C.1.
Computational Cost of Evaluation
Unlike existing EM evaluation, our proposed RAEE relies on advanced LLMs to evaluate the quality of predicted events, leading to much higher costs. The average cost of evaluating one model on a dataset is between 2-5\$, which is smaller than established benchmarks like AlpacaEval Li et al. (2023b).
Insufficient Evaluation for LLMs
@@
Appendix A Datasets Statistics
We utilize 10 widely-used datasets as evaluation benchmarks in both EM and RAEE evaluation. Since these datasets originate from various domains and differ in terms of pre-defined ontologies as well as size, it is necessary to introduce basic statistics of these datasets, as shown in Table 8.
@@ Appendix A Datasets Statistics
Appendix B Key Criteria in RAEE
Considering that the requirements for EE vary across different scenarios and applications, we introduce key criteria in RAEE to enable the adaptive mechanism. As shown in Tabel 9, these key criteria are based on factors that significantly influence the evaluation results in our human annotation meta-evaluation experiment. Furthermore, this adaptive mechanism is easily extendable, allowing for the addition of specific judgment criteria to account for more factors in extraction.
@@ Appendix B Key Criteria in RAEE
Appendix C Prompts
C.1 Reassessment Prompts in RAEE
@@
C.2 Inference Prompts for LLMs Extracting Events
Table 12 and 14 illustrate the prompts we use for testing the event extraction ability of LLMs in ED and EAE tasks, respectively. It should be noted that we follow existing closed domain EE studies Gao et al. (2023); Wei et al. (2024); cai-etal-2024-improving-event to explicitly provide all possible types in the inference prompt. This is a necessary step when performing EE inference with LLMs and does not constitute information leakage.
@@ C.2 Inference Prompts for LLMs Extracting Events
Appendix D Implementation Details